# Perverse microsheaves

### v2

@@ 7.2. The perverse t-structure on constructible sheaves
We wish to apply Lemma 7.2. We should therefore show that the images of the truncation functors, when applied to objects $F\in sh_{\Lambda}(M)$, remain in $sh_{\Lambda}(M)$. Suppose not; consider any smooth Lagrangian point in $ss(\tau_{\leq 0}F)\setminus ss(F)$. (There must be such a point by the assumption that the microsupport is complex analytic Lagrangian.) But by construction, the truncation functors commute with $\omega_{p}$, whence it follows that $\omega_{p}^{-1}\tau_{\leq 0}F[-2r(p)]=\tau_{\leq 0}\omega_{p}^{-1}F[-2r(p)]=0$. A contradiction. The conclusion now follows from Lemma 7.2.
We now review the results of Waschkies [27]. Consider $T^{\circ}M=T^{*}M\setminus M$, and the map $\pi:T^{\circ}M\to\mathbb{P}^{*}M$.
Lemma 7.9.
The natural map
(19) $\mu sh^{pre}(\pi^{-1}(m))\to(\mathbb{P}\mu sh)_{m}$
is an equivalence.
@@
We first check fully-faithfulness. It is equivalent to check that the natural map
$\mu sh^{pre}(\pi^{-1}(m))\to(\pi_{*}\mu sh)_{m}$
is fully-faithful. Given objects $K,L$ in the source, Waschkies [27, Sec. 2.4] shows that $\hom(K,L)$ in the source is computed by $(\pi_{*}\mu\operatorname{Hom}(K,L))_{m}$, which also computes $\hom(K,L)$ in the target.
We now check essential surjectivity. If $K\in(\mathbb{P}\mu sh)_{m}$, then it is represented by an object $\tilde{K}\in\mathbb{P}\mu sh(U),p\in U,$ which is representable by a sheaf at $p$. This means that (after possibly shrinking $U$) there is a complex-contact embedding $\chi:U\to V\subset\mathbb{P}^{*}N$ so that $\chi_{*}(K)=\chi_{*}(\tilde{K})\in(\pi_{*}\mu sh)_{\chi(m)}$ is in the image of $(\pi_{*}\mu sh^{pre})_{\chi(m)}\to(\pi_{*}\mu sh)_{\chi(m)}$. By standard quantization of sheaf kernels (see e.g. [27, Thm. 4.2.1]) we have
@@
where the vertical arrows are (19). The claim follows. ∎
Theorem 7.10.
[27, Thm. 5.1.5] Assume $M$ is complex analytic, $\Lambda\subset T^{\circ}M$ is a $\mathbb{C}^{*}$ invariant (singular) Lagrangian, and $\Lambda/\mathbb{C}^{*}\to M$ is a finite map. Then the natural map of sheaves of categories
$sh_{\Lambda}/loc\to\mathbb{P}\mu sh_{\Lambda}$
is an equivalence.
@@
We check at stalks. There are natural maps:
$(sh_{\Lambda}/loc)_{m}=sh_{\Lambda,m}/loc_{m}\to\mu sh^{pre}(\pi^{-1}(\tilde{m}))\to(\mathbb{P}\mu sh_{\Lambda})_{\tilde{m}}$
where $m\in M$ and $\tilde{m}\in\Lambda$ its (possibly disconnected) preimage. In [27, Thm. 5.1.5], Waschkies constructs the inverse using the refined microlocal cutoff of D’Agnolo [6], and shows that it is fully faithful essentially by arguing that all categories in question have homs given by the $\mu hom$ functor, similarly to [15, Thm. 6.1.2]. ∎
Remark 7.11.
Note that one must sheafify quotients, so $(sh_{\Lambda}/loc)(U)$ is generally not $sh_{\Lambda}(U)/loc(U)$. However, both $sh_{\Lambda}$ and $loc$ are constructible, so for small enough open neighborhood $U$ of any $m\in M$, we have $(sh_{\Lambda}/loc)(U)=(sh_{\Lambda}/loc)_{m}=(sh_{\Lambda})_{m}/loc_{m}=sh_{\Lambda}(U)/loc(U)$.
Remark 7.12.
The analogue of Theorem 7.10 with complex analyticity hypotheses removed and with $\mathbb{C}^{*}$ replaced by $\mathbb{R}^{+}$ is also true (indeed easier: the cutoffs already in [15] are good enough and one does not need [6]); see [22, Lem 6.7, Prop. 6.9].
Corollary 7.13.
In the situation of Theorem 7.10, there is a $t$-structure on $\pi_{*}\mu sh_{\Lambda}$ characterized by the formulas in Definition 7.7.
@@
For any $U\subset M$, the $t$-structure descends to $sh_{\Lambda}(U)/loc(U)$ by Lemma 7.2 and Lemma 7.8. For $p\in T^{\circ}M$, the functor $\omega_{p}$ descends to $sh_{\Lambda}(U)/loc(U)$, and clearly the quotient $t$-structure on $sh_{\Lambda}(U)/loc(U)$ is characterized exactly as in Definition 7.7. ∎
Remark 7.14.
In [27, Thm. 6.1.5], this is extended to arbitrary $\Lambda\subset\mathbb{P}\mathbb{T}^{*}M$ by noting that complex contact transformation preserves the conditions of Definition 7.7 above, using the degree shift formula [15, prop. 7.4.6]. In our discussion below, we don’t explicitly use this complex contact invariance of Definition 7.7 to construct the microlocal perverse $t$-structure, but would need it to check that on cotangent bundles, we have defined the usual perverse $t$-structure.
Definition 7.15.
Fix a $t$-structure on the coefficient category $C$, e.g. the standard $t$-structure on the category of $R$-modules, for a commutative ring $R$. Let $V$ be a contact manifold of complex dimension $12n-$ 1. Fix any $(C,C^{\leq 0})$ orientation data $o$. (If $C=R-mod$, we may always choose the canonical orientation data.)
We define subcategories of $\mu sh_{\tilde{V},\mathbb{C}-c,o}$ by constraining the microstalks to respect the $t$-structure of $C$. In the notation of Definition 5.11:
$(\mu sh_{\tilde{V},\mathbb{C}-c,o})^{\leq 0}:=(\mu sh_{\tilde{V},\mathbb{C}-c,o})^{C^{\leq 0}[2n]}$
$(\mu sh_{\tilde{V},\mathbb{C}-c,o})^{\geq 0}:=(\mu sh_{\tilde{V},\mathbb{C}-c,o})^{C^{\geq 0}[2n]}$
The $[2n]$ is just a convention, set to match the usual conventions for perverse sheaves.
Theorem 7.16.
$(\mathbb{P}\mu sh_{V,\mathbb{C}-c,o})^{\leq 0}$, $(\mathbb{P}\mu sh_{V,\mathbb{C}-c,o})^{\geq 0}$ determine a $t$-structure on $\mathbb{P}\mu sh_{V,\mathbb{C}-c,o}$.

### v3

@@ 6.3. Perverse microsheaves on complex contact and symplectic manifolds
It is sufficient to show that, for $x\in\operatorname{Pic}(\mathcal{C})$, tensoring $x\otimes(-)$ fixes $\mathcal{C}^{\leq 0}$ if and only if tensoring its inverse $x^{-1}\otimes(-)$ fixes $\mathcal{C}^{\geq 0}$. We show the “if” direction, since the argument is symmetric. As remarked in [30, Remark 1.2.1.3], an object $a\in\mathcal{C}$ is in $\mathcal{C}^{\geq 1}$ if $\operatorname{Hom}(b,a)=0$ for all $b\in\mathcal{C}^{\leq 0}$. Thus we consider such objects and compute that $\operatorname{Hom}(b,x^{-1}\otimes a)=\operatorname{Hom}(x\otimes b,a)=0$. This implies that $x^{-1}\otimes(-)$ fixes $\mathcal{C}^{\geq 1}$ but, since $[n]$ commutes with tensor, it fixes $\mathcal{C}^{\geq n}$ for all $n\in\mathbb{Z}$. ∎
Since the collection of subcategories $\{\mathcal{C}^{\leq 0},\mathcal{C}^{\geq 0}\}$ is anchored, there is a notion of constrained gradings and orientations defined in Definition 4.34. As explained by Equation 17, the canonical grading from the complexes structure induces a $(\mathcal{C},\{\mathcal{C}^{\leq 0},\mathcal{C}^{\geq 0}\})$-grading.
Definition 6.15.
Let $o$ be a $(\mathcal{C},\{\mathcal{C}^{\leq 0},\mathcal{C}^{\geq 0}\})$-orientation, i.e., a lifting of the induced grading to a $\mathcal{C}$-Maslov data. We denote by $\mu sh_{\widetilde{V},o}$ the associated sheaf of microsheaves on $\widetilde{V}$ and similar notations for the subsheaves with support conditions.
Let $L$ be a complex Legendrian. Then Lemma 3.12 allows us to choose constrained secondary Maslov data, in the sense defined above Corollary 4.36, and the cited Proposition implies that, for any constructible microsheaf $\mathcal{F}\in\mu sh_{\widetilde{V},\mathbb{C}-c,o}$ and any smooth point $p\in\operatorname{supp}(\mathcal{F})^{\circ}$, whether the microstalk $\omega_{p}^{-1}\mathcal{F}$ is in $\mathcal{C}^{\leq 0}$ or $\mathcal{C}^{\geq 0}$ (and hence all their shiftings) is a well-defined notion. Thus, we have the following generalization of Definition 6.12:
Definition 6.16.
Fix a $t$-structure on the coefficient category $\mathcal{C}$. Let $V$ be a contact manifold of complex dimension 2n12n-1. Fix any $(\mathcal{C},\{\mathcal{C}^{\leq 0},\mathcal{C}^{\geq 0}\})$-orientation data $o$. We define subcategories of $\mu sh_{\widetilde{V},\mathbb{C}-c,o}$ by constraining the microstalks to respect the $t$-structure of $\mathcal{C}$:
$(\mu sh_{\widetilde{V},\mathbb{C}-c,o})^{\leq 0}:=\{\mathcal{F}\in\mu sh_{V,\mathbb{C}-c,o}\,|\,p\in\operatorname{supp}(\mathcal{F})^{\circ}\implies\omega_{p}^{-1}\mathcal{F}[-n]\in\mathcal{C}^{\leq 0}\}$
$(\mu sh_{\widetilde{V},\mathbb{C}-c,o})^{\geq 0}:=\{\mathcal{F}\in\mu sh_{V,\mathbb{C}-c,o}\,|\,p\in\operatorname{supp}(\mathcal{F})^{\circ}\implies\omega_{p}^{-1}\mathcal{F}[-n]\in\mathcal{C}^{\geq 0}\}$
As perversity can be checked locally, the exact same argument of Theorem 6.17 and Theorem 1.6 implies the following theorem.
Theorem 6.17.
Fix a $t$-structure on the coefficient category $\mathcal{C}$. Let $V$ be a contact manifold of complex dimension 2n12n-1. Or, similarly, let $W$ be an exact complex symplectic manifold of complex dimension $2n$ with a $\mathbb{C}^{*}$-action of weight $1$.
Given any $(\mathcal{C},\{\mathcal{C}^{\leq 0},\mathcal{C}^{\geq 0}\})$-orientation data $o$ on $V$ (resp. $W$), the pairs
$\left((\mu sh_{V,\mathbb{C}-c,o}\right)^{\leq 0},(\mu sh_{V,\mathbb{C}-c,o})^{\geq 0})\left(\text{resp.}\ \left(\left((\gamma_{\mathbb{C}})_{*}\mu sh_{W,\mathbb{C}-c}\right)^{\geq 0}\right),\left((\gamma_{\mathbb{C}})_{*}\mu sh_{W,\mathbb{C}-c}\right)^{\leq 0}\right)$
determine a $t$-structure on $\mu sh_{V,\mathbb{C}-c,o}$ (resp. $(\gamma_{\mathbb{C}})_{*}\mu sh_{W,\mathbb{C}-c}$). Furthermore, for $\mathcal{F},\mathcal{G}\in\mu sh_{V,\mathbb{C}-c,o}^{\heartsuit}$ $\left(resp.\left((\gamma_{\mathbb{C}})_{*}\mu sh_{W,\mathbb{C}-c})^{\heartsuit}\right)\right)$, the sheaf of morphisms
$\mathcal{H}om_{\mu sh_{V}}(\mathcal{G},\mathcal{F})[\frac{1}{2}\dim\widetilde{V}]\left(\text{resp.}\ \mathcal{H}om_{(\gamma_{\mathbb{C}})_{*}\mu sh_{W}}(\mathcal{G},\mathcal{F})[\frac{1}{2}\dim W]\right)$
is a perverse sheaf on the symplectization $\widetilde{V}$ (resp. $W$).
Remark 6.18.
A $t$-structure is said to be nondegenerate if $\bigcap\mathcal{C}^{\leq 0}=0=\bigcap\mathcal{C}^{\geq 0}$. By co-isotropicity of microsupport, the vanishing of all microstalks implies the vanishing of an object. We conclude that if the $t$-structure on $\mathcal{C}$ is non-degenerate, then the $t$-structure on $\mu sh_{V,\mathbb{C}-c,o}$ is also non-degenerate.
Remark 6.19.
Choose a stable (not necessarily presentable) subcategory $\mathcal{D}\subset\mathcal{C}$ to which the $t$-structure restricts. Require $(\mathcal{C},\{\mathcal{C}^{\leq 0},\mathcal{C}^{\geq 0},\mathcal{D}\})$-orientation data. Then it is evident from the definitions that the truncation functors preserve, hence define a $t$-structure on, the subcategory of objects with all microstalks in $\mathcal{D}$, characterized by the same formulas save only with e.g. $\mathcal{C}^{\leq 0}$ replaced by $\mathcal{C}^{\leq 0}\cap\mathcal{D}$. For instance, we can take various bounded categories e.g. $\mathcal{D}=\mathcal{C}^{+}=\bigcup\mathcal{C}^{\geq n}$, or $\mathcal{D}=\mathcal{C}^{-}=\bigcap\mathcal{C}^{\leq n}$, or $\mathcal{D}=\mathcal{C}^{b}=\mathcal{C}^{+}\cap\mathcal{C}^{-}$, or ask the microstalks to be compact objects $\mathcal{D}=\mathcal{C}^{c}$.
Remark 6.20.
Fix $D\subset C$ as above, and assume $D^{\heartsuit}$ is Artinian. (E.g., C=kmodC=k-mod for a field $k$ and $D=C^{c}$.) Then the full subcategory of $(\mu sh_{V,\mathbb{C}-c,o})^{D,\heartsuit}$ on objects with finitely stratified (rather than just locally finite) support is also Artinian. Indeed, any descending chain must have eventually stabilizing microsupports; we may restrict attention to one each in the finitely many connected components of the smooth locus of the support, hence by some point, all have stabilized.
Remark 6.21.
Let $V$ be a complex contact manifold and let $L\subset V$ be a complex Legendrian. Then $\mu sh_{V,L}(-)$ is a sheaf of stable categories while $\mathbb{P}erv_{V,L}(-)$ is a sheaf of abelian categories. Be warned however that $D(\mathbb{P}erv_{V,L}(-))$ is only a presheaf of stable categories. In particular, the natural map $D(\mathbb{P}erv_{V,L}(-))\to\mu sh_{V,L}(-)$ may restrict to an equivalence on stalks without being an equivalence on global sections. A very special case: let $V=T^{*}S^{2}\times\mathbb{\mathbb{C}}$ and $L=0_{S^{2}}\times\{0\}$. Then $\mu sh_{V,L}(L)=loc(S^{2})\otimes loc(\mathbb{C}^{*})$ while $\mathbb{P}erv_{V,L}(L)=vect_{\mathbb{C}}\otimes loc(\mathbb{C}^{*})$, due to $S^{2}$ being simply connected. Similarly, $\mu sh_{T^{*}\mathbb{S}^{2},0_{S^{2}}}(0_{S^{2}})=loc(S^{2})$ while $\mu sh_{T^{*}\mathbb{S}^{2},0_{S^{2}}}(0_{S^{2}})^{\heartsuit}=vect_{\mathbb{C}}$.
Appendix A Existence of orientation data for real symplectic manifolds
            by Sanath Devalapurkar
We keep the notation of Section 3. We write $\operatorname{Sq}^{2}:B(\mathbb{Z}/2)\to B^{3}(\mathbb{Z}/2)$ for the map between infinite loop spaces representing the second Steenrod operation $\operatorname{Sq}^{2}:H^{*}(-;\mathbb{Z}/2)\rightarrow H^{*+2}(-;\mathbb{Z}/2)$.
Lemma A.1 ( [9, (2.3)]).
The connecting morphism for the exact triangle
$B^{2}(\mathbb{Z}/2)\to\tau_{\leq 2}(U/O)\to\tau_{\leq 1}(U/O)=B\mathbb{Z}\to$
is the composition $B\mathbb{Z}\to B(\mathbb{Z}/2)\xrightarrow{\operatorname{Sq}^{2}}B^{3}(\mathbb{Z}/2)$. ∎
We consider the following diagram:
[…]
The map from $U(1)\to B\sqrt{SU}$ is the composition $U(1)\to BSU\to B\sqrt{SU}$, where the first map is the connecting map of the fiber sequence $SU\to U\xrightarrow{det}U(1)$. The bottom row is a fiber sequence, so the map $\alpha$ is induced by the fact that the composition $B\sqrt{SU}\to\pi_{\leq 2}B(U/O)\simeq B^{2}\mathbb{Z}$ is null.
To establish the existence of the dotted arrow $U(1)\to B\mathbb{Z}$, it is enough to prove that the composition $B(SU)\to B(\sqrt{SU})\to B^{3}(\mathbb{Z}/2)$ is null (since $U(1)\to B\sqrt{SU}$ factors through $B(SU)\to B(\sqrt{SU})$). To this end, consider the fiber sequence $SU\to U\to U(1)$. Since $U(1)=\tau_{\leq 2}U$, it follows that $SU=\tau_{\geq 3}U$. Hence $B(SU)$ is $3$-connected, so any map into $B^{3}(\mathbb{Z}/2)$ is null.
Lemma A.2.
The map $\alpha:B\sqrt{SU}\rightarrow B^{3}(\mathbb{Z}/2)$ factors as a map of infinite loop spaces
$B\sqrt{SU}\xrightarrow{B\det}B(\mathbb{Z}/2)\xrightarrow{\operatorname{Sq}^{2}}B^{3}(\mathbb{Z}/2).$
@@ Appendix A Existence of orientation data for real symplectic manifolds
This follows by contemplating the diagram:
@@
The commutativity of the leftmost square comes from the fiber sequences
$U{U}U(1){{U(1)}}BSU{{B\sqrt{SU}}}$$\scriptstyle{\operatorname{det}}$$scriptstyle{z\mapsto z^{2}}$$\scriptstyle{\operatorname{det}^{2}}$
and the rightmost square comes from (25). It remains to explain the commutativity of the bottom triangle. Since $B(SU)\to B(\sqrt{SU})\to B(\mathbb{Z}/2)$ is a fiber sequence, it is enough to prove that the composition $B(SU)\to B(\sqrt{SU})\to B^{3}(\mathbb{Z}/2)$ is null, which was proved above. ∎
Proposition A.3.
A choice of null-homotopy of $\operatorname{Sq}^{2}:B\mathbb{Z}/2\to B^{3}\mathbb{Z}/2$ gives a section of the forgetful map from grading/orientation data to grading data.
@@
A grading is a lift of the natural map $X\to BU$ to some $f:X\to B\sqrt{SU}$. Giving orientation data is giving a null-homotopy of the map $\alpha\circ f$, which by Lemma A.2, can be written as $\operatorname{Sq}^{2}\circ B\det\circ f$. ∎
While $\operatorname{Sq}^{2}$ is nonzero as a map of infinite loop spaces (as the Steenrod square is a nontrivial operation on cohomology), it is null as a map of spaces. Indeed, under the identifications
(26) $\pi_{0}\left(\operatorname{Map}(B(\mathbb{Z}/2),B^{3}(\mathbb{Z}/2)\right)=H^{3}(B(\mathbb{Z}/2);\mathbb{Z}/2)=H^{3}(\mathbb{R}\mathbb{P}^{\infty},\mathbb{Z}/2)=\mathbb{Z}/2\langle w_{1}^{3}\rangle,$
the (homotopy class of the) map $\operatorname{Sq}^{2}:B(\mathbb{Z}/2)\rightarrow B^{3}(\mathbb{Z}/2)$ corresponds to the element $\operatorname{Sq}^{2}(w_{1})\in\mathbb{Z}/2\langle w_{1}^{3}\rangle$. But $\operatorname{Sq}^{2}(w_{1})=0$, since Steenrod squares have the well-known property that $\operatorname{Sq}^{n}(x)=0$ if $n>\deg(x)$.
Now, the homotopy classes of null-homotopies of $\operatorname{Sq}^{2}$ are given by $B(\mathbb{Z}/2),B^{2}(\mathbb{Z}/2)]=\mathrm{H}^{2}(\mathbb{R}\mathbb{P}^{\infty},\mathbb{Z}/2)=\mathbb{Z}/2$. Consider the inclusion $O\to\sqrt{SU}$; to study secondary orientation data, we will be interested in null-homotopies of the composition $\alpha^{\prime}:BO\to B\sqrt{SU}\xrightarrow{\alpha}B^{3}(\mathbb{Z}/2\mathbb{Z})$. The space of such null-homotopies is a torsor for $\operatorname{Map}(BO,B^{2}(\mathbb{Z}/2\mathbb{Z}))$, the homotopy classes of which are $H^{2}(BO,\mathbb{Z}/2\mathbb{Z})=\mathbb{Z}/2\langle w_{1}^{2},w_{2}\rangle$.
One such null-homotopy arises from the null-homotopy of the factorization $O\to\sqrt{SU}\to\sqrt{SU}/O=B^{2}O\xrightarrow{Bw_{1}}B^{2}(\mathbb{Z}/2\mathbb{Z})$; let us denote this null-homotopy by $\tau$. Two more such null-homotopies arise by noticing that $\alpha^{\prime}$ factors through $BO\to B\sqrt{SU}\to B(\mathbb{Z}/2)\xrightarrow{\operatorname{Sq}^{2}}B^{3}\mathbb{Z}/2$, and then composing with one of the two null-homotopies of $\operatorname{Sq}^{2}$.
Lemma A.4.
The two null-homotopies of $\alpha^{\prime}:BO\to B^{3}(\mathbb{Z}/2\mathbb{Z})$ induced by composition with null-homotopies of $\operatorname{Sq}^{2}$ represent the classes $\tau+w_{2}$ and $\tau+w_{2}+w_{1}^{2}$.
@@
Denote by $\bar{\tau}$ the null-homotopy $BU\rightarrow B^{3}(\mathbb{Z}/2\mathbb{Z})$ by pre-composing $\tau$ with $BU\rightarrow BO$. As the pullback map $H^{2}(BO,\mathbb{Z}/2)\rightarrow H^{2}(BU,\mathbb{Z}/2)$ is given by $w_{2}\mapsto c_{1}$ and $w_{1}^{2}\mapsto 0$, it is sufficient to show that the two null-homotopies both go to $\bar{\tau}+c_{1}$ after further composing to $BU\rightarrow BO\rightarrow B^{3}(\mathbb{Z}/2\mathbb{Z})$. Thus, we consider the composition $U\rightarrow O\to\sqrt{SU}\to\sqrt{SU}/O$. The commutative diagram right above Definition 3.1 provides the following:
[…]
We note that the dashed map exists, since any map of the form $BSp\rightarrow B^{3}(\mathbb{Z}/2\mathbb{Z})$ factorizes to $BSp\rightarrow\tau_{\leq 3}BSp\rightarrow B^{3}(\mathbb{Z}/2\mathbb{Z})$ which is canonically homotopic as $\tau_{\leq 3}BSp=0$. In particular, the two homotopies induced from null-homotopy of $\alpha$ becomes the same after composing with $BU\rightarrow BO$ and is given by the fiber sequence $BSp\rightarrow B(Sp/U)\rightarrow B^{2}U$. On the other hand, the left three horizontal maps form a map between fiber sequences. Thus, $\bar{\tau}$ comes from post-composing the fiber sequence $BU\rightarrow BSp\rightarrow B(Sp/U)$. By the proof of Proposition 3.10, we see that their difference is given by $BU\xrightarrow{c_{2}}B^{2}(\mathbb{Z})\rightarrow B^{2}(\mathbb{Z}/2\mathbb{Z})$, which is exactly what we want. ∎
We write $\nu_{+}$ for the null-homotopy with class $\tau+w_{2}$ and $\nu_{-}$ for the one with class $\tau+w_{2}+w_{1}^{2}$.
Lemma A.5.
If $\nu_{\pm}$ is used to define orientation data on a symplectic manifold $X$, then secondary orientation data on a Lagrangian $L$ is a $Pin_{\pm}$ structure on $L$.